After volleyball practice your team of 15 players usually eats 36 Krazy Kreme donuts. Using the same ratio, how many donuts should you plan on buying for upcoming tournament which has 40 players? 

Group of answer choices




1
Expert's answer
2021-11-08T16:01:18-0500

Let p= number of players

and d= number of donuts.


"p\propto d"

"\implies p=kd\quad where" "k"

is a constant of proportion.


"When\> p=15\quad d=36"

"\therefore15= k.36\quad\implies k= \frac{15}{36}"


When "p=40\quad 40= \frac{15}{36}d"


"\frac{40*36}{15}= 96"


"96" donuts